# Trajectory Balance with Asynchrony: Decoupling Exploration and Learning for Fast, Scalable LLM Post-Training

### v1

@@ Abstract
Reinforcement learning (RL) is a critical component of large language model (LLM) post-training. However, existing on-policy algorithms used for post-training are inherently incompatible with the use of experience replay buffers, which can be populated scalably by distributed off-policy actors to enhance exploration as compute increases. We propose efficiently obtaining this benefit of replay buffers via Trajectory Balance with Asynchrony (TBA), a massively scalable LLM RL system. In contrast to existing approaches, TBA uses a larger fraction of compute on search, constantly generating off-policy data for a central replay buffer. A training node simultaneously samples data from this buffer based on reward or recency to update the policy using Trajectory Balance (TB), a diversity-seeking RL objective introduced for GFlowNets. TBA offers three key advantages: (1) decoupled training and search, speeding up training wall-clock time by 4x or more; (2) improved diversity through large-scale off-policy sampling; and (3) scalable search for sparse reward settings. On mathematical reasoning, preference-tuning, and automated red-teaming (diverse and representative post-training tasks), TBA produces speed and performance improvements over strong baselines.
Figure 1: TBA performs rapid, scalable exploration of model responses, improving RL efficiency on the GSM8K mathematical reasoning task. All plotted points use 4xA100 GPUs (or comparable L40S GPUs). DPO and RLOO baselines taken from Noukhovitch et al. (2025), PPO and VinePPO baselines taken from Kazemnejad et al. (2024). The baseline model is the SFTed RhoMath-1B (Lin et al., 2024) model, which obtains 40.3% accuracy after SFT and before RL. Appendix A has details.
@@ 1 Introduction
Post-training through reinforcement learning (RL) is a critical step in enhancing large language models (LLMs), aligning them with human preferences and improving their reasoning abilities (Christiano et al., 2017). However, widely used RL algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and REINFORCE Leave-One-Out (RLOO) (Ahmadian et al., 2024) suffer from a fundamental limitation: they are on-policy, meaning that data generation and policy updates occur sequentially. This dependence creates bottlenecks that reduce resource utilization. Moreover, the benefit of scaling on-policy data generation may be limited (Hou et al., 2024).
We introduce Trajectory Balance with Asynchrony (TBA), a distributed RL framework designed to efficiently and scalably leverage compute for LLM post-training. TBA decouples data generation from policy updates using an off-policy training objective based on Trajectory Balance (TB) (Malkin et al., 2022a). The framework uses multiple searcher nodes that independently generate diverse trajectories and store them in a central replay buffer, while a single trainer node asynchronously samples from this buffer to update the policy. By removing the dependency between data generation and training, TBA ensures high resource utilization and facilitates scalable search. By computing model updates with large, efficiently-generated off-policy response sets, the TB objective leverages response scaling to produce better models than popular RL baseline methods.
TBA offers three key advantages over existing RL-based LLM post-training approaches: 1) Asynchrony enables massive parallelization, significantly reducing training wall-clock time as shown in Figures 1 and 3. 2) Diverse off-policy sampling from a replay buffer improves exploration and prevents mode collapse. 3) Scalable search capabilities make TBA particularly effective in sparse reward settings, such as automated red-teaming.
We validate TBA on mathematical reasoning, preference-tuning, and automated red-teaming. Our results show that TBA achieves performance comparable to or better than existing methods while significantly improving training speed. Our key contributions are summarized as follows:
  • We introduce TBA, a novel distributed RL framework for LLM post-training.
  • We decouple data generation and policy updates, improving training speed and scalability.
  • We demonstrate the effectiveness of trajectory balance for LLM post-training, exploiting its ability to efficiently leverage large-scale off-policy data.
  • We demonstrate significant speedups (4x or more) in RL for mathematical reasoning, preference-tuning, and automated red-teaming.
By enabling high-quality and fast off-policy post-training, TBA contributes to the broader goal of scalable and effective LLM alignment, ensuring that large models can be refined more efficiently for real-world deployment.
@@ 2 Related Work
Reinforcement Learning (RL) has been an integral component for training LLMs (Google Gemini Team, 2024; OpenAI, 2023). In particular, RL has became the de facto approach for aligning language models with human preferences Christiano et al. (2017); Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022). Much of this work relies on Proximal Policy Optimization (PPO) (Schulman et al., 2017), an on-policy RL algorithm which has become a default choice for fine-tuning LLMs due to it’s strong performance across different setups. Aside from PPO, other on-policy objectives such as REINFORCE (Ahmadian et al., 2024) and variants like GRPO (Shao et al., 2024) and VinePPO (Kazemnejad et al., 2024) have also been studied in the context of language models.
An alternative to PPO-based fine-tuning is rejection sampling fine-tuning, inspired by the best-of-$n$ inference approach proposed by Nakano et al. (2021). Recent work by Dong et al. (2023); Gulcehre et al. (2023), and Wang et al. (2024) extends this concept by generating $n$ candidate responses for each prompt, ranking them with a learned reward function, and fine-tuning the model based on the highest-ranking responses. On the other hand, direct preference learning approaches (Rafailov et al., 2023; Azar et al., 2024; Tang et al., 2024) skip reward modeling entirely and train language models to directly optimize responses under a preference model. Finally, Hu et al. (2024) introduced GFlowNet fine-tuning, leveraging off-policy GFlowNet algorithms for fine-tuning language models, which we build upon in this work.
@@
Distributed RL spreads actors/searchers, learners, and environments across a collection of computing resources. Asynchronous distributed RL does this such that searcher and trainer processes do not necessarily share the same weights, which can significantly improve training speed (Nair et al., 2015; Wang et al., 2025) and facilitate RL in complex, high-dimensional domains (Hessel et al., 2021; Huang et al., 2023; Horgan et al., 2018).
A foundational method in this area is Asynchronous Advantage Actor-Critic (A3C) (Mnih, 2016). In A3C, multiple parallel workers asynchronously interact with the environment and communicate gradients to a central node. Our approach to async distributed RL more closely resembles the Importance-Weighted Actor-Learner Architecture (IMPALA) method (Espeholt et al., 2018), which communicates experience trajectories (state, action, and reward tuples) to the central node.
@@
Through adversarial interactions, LLM red-teaming clarifies the robustness and risks of a target LLM. Automating the generation of these adversarial scenarios, automated red-teaming frameworks can help uncover vulnerabilities, biases, and unintended behaviors in models more quickly, enabling preemptive mitigation before deployment (Wei et al., 2023; Ji et al., 2024).
Perez et al. (2022) proposed training language models using RL to discover prompts that elicit harmful responses from some target LLM. Standard RL approaches, however, are susceptible to mode collapse and fail to achieve the attack diversity that is critical for successful red-teaming. Hong et al. (2024) introduced a curiosity bonus to encourage generation of diverse red-teaming prompts, whereas Samvelyan et al. (2024) proposed sampling an attack prompt from a pool and iteratively mutating the prompt with auxiliary LLMs. Lee et al. (2025) proposed using GFlowNet fine-tuning followed by MLE smoothing to generate diverse, transferable, and effective prompts. Our red-teaming experiments augment their TB objective optimization with our distributed asynchronous framework.
@@ 3 Preliminaries
KL regularized RL as probabilistic inference
We study the problem of fine-tuning a pretrained language model $\pi_{\text{ref}}$ with a reward model $r_{\phi}$. For mathematical reasoning (Cobbe et al., 2021) our reward model simply computes a response’s correctness, while our preference-tuning for alignment (Ziegler et al., 2019) and red teaming (Perez et al., 2022) experiments use reward models optimized with human preference data. Notably, reward maximizing RL with learned reward functions is susceptible to spurious modes of the reward (Skalse et al., 2022; Pan et al., 2022; Gao et al., 2023), resulting in poor performance and low diversity in responses. This is addressed by constraining the fine-tuned model to be close to the initial model in terms of the KL divergence:
$\begin{split}\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\mathbb{E}_{\mathbf{y}\sim\pi(\mathbf{y}\mid\mathbf{x})}[r_{\phi}(\mathbf{y};\mathbf{x})]\\ -\beta\mathbb{D}_{\text{KL}}(\pi(.\mid\mathbf{x})||\pi_{\text{ref}}(.\mid\mathbf{x}))].\end{split}$ (1)
Online RL algorithms such as PPO (Schulman et al., 2017; Stiennon et al., 2020) and REINFORCE (Williams, 1992; Kool et al., 2019; Ahmadian et al., 2024) can be used to optimize the model (which is the policy), whereas offline objectives such as DPO (Rafailov et al., 2023) can be used to train the model directly on the preference data and achieve the same optimal policy asymptotically.
Eq. 1 can be interpreted from a probabilistic perspective as a Bayesian posterior inference problem (Korbak et al., 2022). The optimal policy for Eq. 1 is given as:
@@
Approaches such as Gibbs sampling can be used to produce samples from the optimal policy without any fine-tuning (Xiong et al., 2024). These MCMC approaches can be very expensive at inference time and intractable for long sequences. Within the probabilistic interpretation, on-policy RL to maximize Eq. 1 is equivalent to amortized variational inference to minimize the reverse KL with respect to the posterior density (Korbak et al., 2022). However, reverse KL optimization is susceptible to mode collapse and requires on-policy samples. The reliance on on-policy samples can limit the scalabality as we cannot use replay buffers that can be populated in parallel at scale, since new updates of the policy invalidate the on-policy nature of the older replay buffer examples. In practice this means that the policy can get stuck in suboptimal solutions and stop exploring, or may obtain high reward at the cost of diversity. This motivates us to consider an alternative off-policy amortized variational inference to efficiently leverage scalable computational resources for flexible exploration.
GFlowNets
Generative Flow Networks (GFlowNets; Bengio et al., 2021; 2023) are a framework for off-policy training of hierarchical generative models to sample proportional to a given unnormalized density (reward) function. GFlowNets frame probabilistic inference as a sequential decision-making problem, learning a policy to construct the objects (e.g. sequences) by putting together building blocks (e.g. tokens) and optimizing consistency-based objectives. GFlowNet objectives have been used for fine-tuning autoregressive (Hu et al., 2024; Lee et al., 2025) as well as discrete diffusion language models (Venkatraman et al., 2024).
To fine-tune a language model to sample from Eq. 2, we can set as a reward $R(\mathbf{y};\mathbf{x})=\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})\exp(\beta^{-1}r_{\phi}(\mathbf{y};\mathbf{x}))$. Following Lee et al. (2025), we use the trajectory balance objective (Malkin et al., 2022b) for training the language model policy $\pi_{\theta}$, which is defined over a response $\mathbf{y}$ as
@@
$Z(\mathbf{x})$ is a positive scalar function of the query $\mathbf{x}$, and the response y is a sequence of tokens. When $\mathcal{L}_{\text{TB}}$ is minimized, $Z(\mathbf{x})$ is the partition function of the posterior (i.e. $Z(\mathbf{x})=\sum_{\mathbf{y}}R(\mathbf{y};\mathbf{x})$). Instead of training a value network for $Z$, we use the VarGrad variant of trajectory balance which replaces a learned $Z$ with a batch estimate  (Richter et al., 2020; Nüsken & Richter, 2021; Zhang et al., 2023; Sendera et al., 2024; Venkatraman et al., 2024).
@@
Given $K$ responses $\{\mathbf{y}^{(i,j)}\}_{j=1}^{K}$ for a query $\mathbf{x}^{(i)}$, a batch estimate of $Z$ can be computed as follows
@@
$\begin{split}\log\hat{Z}(\mathbf{x}^{(i)})=\frac{1}{K}\sum_{j=1}^{K}\bigg{(}\sum_{t=1}^{T}\log\pi_{\text{ref}}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})-\log\pi_{\theta}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})+\frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\bigg{)}.\end{split}$ (4)
The estimate $\hat{Z}$ can be plugged into Eq. 3 for a batch $\mathbf{B}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i,j)})_{j=1}^{K}\}_{i=1}^{B}$ to get
@@
$\begin{split}\mathcal{L}_{\text{TB}}^{\text{VarGrad}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}\Bigg{(}\log\hat{Z}(\mathbf{x}^{(i)})+\sum_{t=1}^{T}\log\pi_{\theta}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})\\ -\log\pi_{\text{ref}}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})-\frac{1}{\beta}r(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\Bigg{)}^{2}.\end{split}$ (5)
@@
An important property of the trajectory balance is that it is off-policy. During training, $\mathbf{y}$ can be sampled from any distribution with full support (Bengio et al., 2021). This enables the use of various exploration strategies (Rector-Brooks et al., 2023; Kim et al., 2024) as well as the use of replay buffers (Shen et al., 2023; Vemgal et al., 2023). In the context of fine-tuning language models, this off-policy nature of the objective makes it a natural choice for large-scale distributed training.

### v2

@@ Abstract
Reinforcement learning (RL) is a critical component of large language model (LLM) post-training. However, on-policy algorithms used for post-training are not naturally robust to a diversified content of experience replay buffers, which asynchronous off-policy actors can efficiently populate in parallel to training. We propose efficiently learning on such off-policy data via Trajectory Balance with Asynchrony (TBA), an approach to asynchronous RL for LLMs that leverages the principled off-policy TB objective. On math, preference-tuning, and automated red-teaming tasks, we post-train models ranging from Pythia 410M to Qwen 2.5 7B, finding TBA offers speed and performance boosts over strong baselines like Online DPO and Dr. GRPO. Beyond TBA’s performance benefits (high accuracy even as asynchrony grows) and speedups ($4\times$ or more), we show its reward- and recency-prioritizing sampling enable further gains as data generation is scaled. Our code is available at https://github.com/bbartoldson/TBA.
Figure 1: TBA excels on the GSM8K mathematical reasoning task. All plotted points use 4xA100 GPUs (or comparable L40S GPUs). DPO and RLOO baselines taken from noukhovitch2024asynchronous, PPO and VinePPO baselines taken from kazemnejad2024vineppounlockingrlpotential. The baseline model is the SFTed RhoMath-1B [lin2024rho] model, which gets 40% accuracy after SFT and before RL. Appendix B has details.
@@ 1 Introduction
Post-training through reinforcement learning (RL) is critical for enhancing large language models (LLMs), aligning them with human preferences and improving their reasoning abilities [christiano2017deep, jaech2024openai]. However, widely used RL algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximal] and REINFORCE Leave-One-Out (RLOO) [ahmadian2024back] suffer from a fundamental limitation: they are on-policy, meaning that data generation and policy updates occur sequentially. This dependence creates bottlenecks that reduce resource utilization. Further, benefits of scaling on-policy data generation may be limited [hou2024does].
We introduce Trajectory Balance with Asynchrony (TBA), an asynchronous RL approach designed to efficiently and scalably leverage compute for LLM post-training. TBA uses an off-policy training objective based on Trajectory Balance (TB) [malkin-etal-2022-balanced], which we hypothesize can improve learning when using async RL to decouple data generation from policy updates, mitigating key RL bottlenecks.
Across mathematical reasoning, preference-tuning, and automated red-teaming tasks, we find TBA broadly offers three advantages over existing LLM post-training approaches: 1) Stable off-policy RL, unlocking asynchrony for massive parallelization and reduced wall-clock times – see Figures 1 and 3. 2) Sampling from a diverse replay buffer, improving exploration and preventing mode collapse. 3) Scalable search compatibility that aids sparse reward settings like automated red-teaming.
Our key contributions are summarized as follows:
  • We introduce TBA, a distributed asynchronous RL framework for fast, scalable LLM post-training.
  • We show that LLMs post-trained with TBA match or exceed performances from existing methods, illustrating TB’s ability to use off-policy data in asynchronous RL.
  • We demonstrate significant speedups ($4\times$ to $50\times$) for RL across mathematical reasoning, preference-tuning, and automated red-teaming.
By enabling high-quality and fast off-policy post-training, TBA contributes to scalable and effective LLM alignment, ensuring that large models can be refined more efficiently for real-world deployment.
@@ 2 Related Work
Reinforcement learning (RL) has been an integral component for training LLMs [geminiteam2024geminifamilyhighlycapable, openai2023gpt4]. In particular, RL has become the de facto approach for aligning language models with human preferences [christiano2017deep, ziegler2019fine, stiennon2020learning, ouyang2022traininglanguagemodelsfollow]. Much of this work relies on Proximal Policy Optimization (PPO) [schulman2017proximal], an on-policy RL algorithm that has become a default choice for fine-tuning LLMs due to its strong performance across different setups. Aside from PPO, other on-policy objectives such as REINFORCE [ahmadian2024back] and variants like GRPO [shao2024deepseekmath] and VinePPO [kazemnejad2024vineppounlockingrlpotential] have also been studied in the context of language models. An alternative to PPO-based fine-tuning is rejection sampling fine-tuning, inspired by the best-of-$n$ inference approach proposed by nakano2021webgpt. Recent work by dong2023raft, gulcehre2023reinforced, and wang2024arithmetic extends this concept by generating $n$ candidate responses for each prompt, ranking them with a learned reward function, and fine-tuning the model based on the highest-ranking responses. Separately, direct preference learning approaches [rafailov2024direct, azar2024general, tang2024generalized] skip reward modeling entirely and train language models to directly optimize responses under a preference model. hu2024amortizing introduced GFlowNet fine-tuning, leveraging off-policy GFlowNet algorithms for fine-tuning language models, which we build upon here.
@@
Distributed RL spreads actors/searchers, learners, and environments across a collection of computing resources. Asynchronous distributed RL does this such that searcher and trainer processes do not necessarily share the same weights, which can significantly improve training speed [nair2015massively, wang2024distrl] and facilitate RL in complex, high-dimensional domains [hessel2021podracer, huang2023cleanba, horgan2018distributed].
A foundational method in this area is Asynchronous Advantage Actor-Critic (A3C) [mnih2016asynchronous]. In A3C, multiple parallel workers asynchronously interact with the environment and communicate gradients to a central node. Our approach to async distributed RL more closely resembles the Importance-Weighted Actor-Learner Architecture (IMPALA) method [espeholt2018impala], which communicates experience trajectories (state, action, and reward tuples) to the central node. As opposed to standard RL environments, where value-based off-policy learning objectives enable principled async training, language models present a distinct challenge since learning value functions is challenging [guo2025deepseek]. Recent work has also leveraged these Async RL frameworks within the context of LLM post-training for training on near on-policy data [noukhovitch2024asynchronous, gehring2024rlef]. In contrast to these recent works, our approach uses a principled off-policy learning objective, allowing training on any off-policy trajectories.
@@
Through adversarial interactions, LLM red-teaming clarifies the robustness and risks of a target LLM. Automating the generation of these adversarial scenarios, automated red-teaming frameworks can help uncover vulnerabilities, biases, and unintended behaviors in models more quickly, enabling preemptive mitigation before deployment [wei2024jailbroken, ji2024beavertails].
perez-etal-2022-red proposed training language models using RL to discover prompts that elicit harmful responses from some target LLM. Standard RL approaches, however, are susceptible to mode collapse and fail to achieve the attack diversity that is critical for successful red-teaming. hong2024curiositydriven introduced a curiosity bonus to encourage generation of diverse red-teaming prompts, whereas rainbow-teaming proposed sampling an attack prompt from a pool and iteratively mutating the prompt with auxiliary LLMs. lee2024learning proposed using GFlowNet fine-tuning followed by MLE smoothing to generate diverse, transferable, and effective prompts. Our red-teaming experiments augment their TB objective optimization with our distributed asynchronous framework.
@@ 3 Preliminaries
3.1 KL regularized RL as probabilistic inference
We study the problem of fine-tuning a pretrained language model $\pi_{\text{ref}}$ with a reward model $r_{\phi}$. For mathematical reasoning [cobbe2021gsm8k] our reward model simply computes a response’s correctness, while our preference-tuning for alignment [ziegler2019fine] and red teaming [perez-etal-2022-red] experiments use reward models optimized with human preference data. Notably, reward maximizing RL with learned reward functions is susceptible to spurious modes of the reward [reward-hacking, reward-hacking-2, gao2023scaling], which can result in poor performance and low diversity in responses. This is addressed by constraining the fine-tuned model to be close to the initial model in terms of the KL divergence:
$\begin{split}\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\mathbb{E}_{\mathbf{y}\sim\pi(\mathbf{y}\mid\mathbf{x})}[r_{\phi}(\mathbf{y};\mathbf{x})]-\beta\mathbb{D}_{\text{KL}}(\pi(.\mid\mathbf{x})||\pi_{\text{ref}}(.\mid\mathbf{x}))].\end{split}$ (1)
Online RL algorithms such as PPO [schulman2017proximal, stiennon2020learning] and REINFORCE [williams1992simple, kool2019buy, ahmadian2024back] can be used to optimize the model (which is the policy), whereas offline objectives such as DPO [rafailov2024direct] can be used to train the model directly on the preference data and achieve the same optimal policy asymptotically.
Eq. 1 can be interpreted from a probabilistic perspective as a Bayesian posterior inference problem [korbak2022rl]. The optimal policy for Eq. 1 is given as:
@@ 3.1 KL regularized RL as probabilistic inference
Approaches such as Gibbs sampling can be used to produce samples from the optimal policy without any fine-tuning [xiong2024iterative]. These MCMC approaches can be very expensive at inference time and intractable for long sequences. Within the probabilistic interpretation, on-policy RL to maximize Eq. 1 is equivalent to amortized variational inference to minimize the reverse KL with respect to the posterior density [korbak2022rl]. However, reverse KL optimization is susceptible to mode collapse and requires on-policy samples. The reliance on on-policy samples can limit the scalability as we cannot use replay buffers that can be populated in parallel at scale, since new updates of the policy invalidate the on-policy nature of the older replay buffer examples. In practice, this means that the policy can get stuck in suboptimal solutions and stop exploring, or may obtain high reward at the cost of diversity. This motivates us to consider an alternative off-policy amortized variational inference to efficiently leverage scalable computational resources for flexible exploration.
3.2 GFlowNets
Generative Flow Networks [GFlowNets; bengio2021flow, bengio2023gflownetfoundations] are a framework for off-policy training of hierarchical generative models to sample proportionally to a given unnormalized density (reward) function. GFlowNets frame probabilistic inference as a sequential decision-making problem, learning a policy to construct the objects (e.g. sequences) by putting together building blocks (e.g. tokens) and optimizing consistency-based objectives. GFlowNet objectives have been used for fine-tuning autoregressive [hu2024amortizing, lee2024learning] as well as discrete diffusion language models [venkatraman2024amortizing]. GFlowNets are MaxEnt RL algorithms [ziebart2010modeling], and notably equivalent to path consistency learning [nachum2017bridging] in sequence generation problems [bengio2021flow].
To fine-tune a language model to sample from Eq. 2, we can set as a reward $R(\mathbf{y};\mathbf{x})=\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})\exp(\beta^{-1}r_{\phi}(\mathbf{y};\mathbf{x}))$. Following lee2024learning, we use the trajectory balance objective [malkin2022trajectory] for training the language model policy $\pi_{\theta}$, which is defined over a response $\mathbf{y}$ as
@@ 3.2 GFlowNets
$Z(\mathbf{x})$ is a positive scalar function of the query $\mathbf{x}$, and the response y is a sequence of tokens. When $\mathcal{L}_{\text{TB}}$ is minimized, $Z(\mathbf{x})$ is the partition function of the posterior (i.e. $Z(\mathbf{x})=\sum_{\mathbf{y}}R(\mathbf{y};\mathbf{x})$). Otherwise it serves as a control variate, reducing the variance of TB gradients. Instead of training a value network for $Z(\textbf{x})$, we use the VarGrad variant of trajectory balance which replaces a learned $Z(\textbf{x})$ with a K-sample batch estimate  [richter2020vargradlowvariancegradientestimator, nusken2021solving, zhang2023robust, sendera2024improvedoffpolicytrainingdiffusion, venkatraman2024amortizing].
@@
Given $K$ responses $\{\mathbf{y}^{(i,j)}\}_{j=1}^{K}$ for a query $\mathbf{x}^{(i)}$, a batch estimate of $Z(\textbf{x}^{(i)})$ can be computed
@@
$\begin{split}\log\hat{Z}(\mathbf{x}^{(i)})=\frac{1}{K}\sum_{j=1}^{K}\bigg(\log\pi_{\text{ref}}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})-\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})+\frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\bigg).\end{split}$ (4)
The (detached) estimate $\hat{Z}$ can be plugged into Eq. 3 for a batch $\mathbf{B}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i,j)})_{j=1}^{K}\}_{i=1}^{B}$ to get
@@
$\begin{split}\mathcal{L}_{\text{TB}}^{\text{VarGrad}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}\bigg(\text{STOP-GRAD}[\log\hat{Z}(\mathbf{x}^{(i)})]+\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})\\ -\log\pi_{\text{ref}}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})-\frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\bigg)^{2}.\end{split}$ (5)
@@
An important property of the trajectory balance is that it is off-policy. During training, $\mathbf{y}$ can be sampled from any distribution with full support [bengio2021flow]. This enables the use of various exploration strategies [rector2023thompson, kim2023local] as well as the use of replay buffers [shen2023towards, vemgal2023empirical]. In the context of fine-tuning language models, this off-policy nature of the objective makes it a natural choice for asynchronous training.
3.3 Differences from other RL objectives
We compare the gradients of the VarGrad TB variant above and Proximal RLOO, an off-policy variant of RLOO with off-policy robustness stemming from a (clipped) importance-sampling (IS) ratio [noukhovitch2024asynchronous]. Please see Appendix D for additional comparisons. With $K$ samples, the gradient for Proximal RLOO for a sample $\mathbf{y}^{(j)}$ is
$\begin{split}\nabla\mathcal{J}_{\text{P-RLOO}}(\theta)=\frac{\pi_{\theta}(\mathbf{y}^{(j)}|\mathbf{x})}{\pi_{ref}(\mathbf{y}^{(j)}|\mathbf{x})}\hat{A}(\mathbf{y}^{(j)}|\mathbf{x})\nabla_{\theta}\log\pi_{\theta}(\mathbf{y}^{(j)}|\mathbf{x}),\end{split}$ (6)
with advantage $\hat{A}(\mathbf{y}^{(j)}|\mathbf{x})=r(\mathbf{y}^{(j)},\mathbf{x})-\frac{1}{K-1}\sum_{i\neq j}r(\mathbf{y}^{(i)},\mathbf{x})$, and IS clipping omitted for brevity.
The gradient for $\mathbf{y}^{(j)}$ using the TB objective $\mathcal{J}_{\text{TB}}$ corresponding to the loss in Eq. 5 is, up to a multiplicative constant, equal to the following (derivation given in Appendix A):
$\begin{split}\nabla\mathcal{J}_{\text{TB}}(\theta)=\hat{A}(\mathbf{y}^{(j)}|\mathbf{x})\nabla\log\pi_{\theta}(\mathbf{y}^{(j)}\mid\mathbf{x}),\end{split}$ (7)
where $\hat{A}(\mathbf{y}^{(j)}|\mathbf{x})=[r(\mathbf{y}^{(j)},\mathbf{x})-\beta\log\frac{\pi_{\theta}(\mathbf{y}^{(j)}|\mathbf{x})}{\pi_{ref}(\mathbf{y}^{(j)}|\mathbf{x})}]-\frac{1}{K}\sum_{i}[r(\mathbf{y}^{(i)},\mathbf{x})-\beta\log\frac{\pi_{\theta}(\mathbf{y}^{(i)}|\mathbf{x})}{\pi_{ref}(\mathbf{y}^{(i)}|\mathbf{x})})]$. That is, for on-policy data $\mathbf{y}$, $\text{TB}^{\text{VarGrad}}$ is equivalent to mean-baseline REINFORCE [williams1992simple] and a KL-divergence-regularized reward [ziegler2019fine]. Note that we use TB on off-policy data, i.e. not sampled from $\pi_{\theta}$, where this equivalence does not hold.
In practice, we find performance on off-policy data is sensitive to the coefficient $\beta$, with larger values tending to promote stability, and smaller values tending to promote accuracy improvements. We find it possible to obtain both benefits via coefficient annealing schedules or resetting of the reference policy [liu2025prorl]. Notably, Kimi K1.5 and K2 were RL-tuned with resetting of the reference policy and an objective that nearly matches TB’s [team2025kimi, team2025kimi2]. However, these Kimi approaches deviate from Eq. 7 by excluding the average log probability ratio from the control variate. Further, their resetting approach causes the denominator in the log probability ratio to always use the trajectory-generating behavior policy as the reference, whereas TBA can reset less frequently, using the original base policy as a reference longer. In Section 6 and Appendix D, we introduce importance sampling to TBA’s objective in Eq. 7 and illustrate it works well with either CISPO IS clipping [chen2025minimax] or IS ratio masking [IcePop2025, zheng2025prosperity].